OpenStudy (anonymous):

Identify the slope and the y-intercept of the equation. y = -3/4 x + 4

jimthompson5910 (jim_thompson5910):

y = -3/4 x + 4 is in the form y = mx+b where m is the slope and (0, b) is the y-intercept. Since m = -3/4, the slope is -3/4 Because b = 4, the y-intercept is (0, 4)